OpenStudy (anonymous):

show the work.... 3x^2-7x=6

OpenStudy (amistre64):

do you know how to "complete the square"?

OpenStudy (amistre64):

first thing we should do is get rid of that "3" that is hanging on in front; we do that by dividing everything by "3" to get: (3/3)x^2 -(7/3)x = (6/3) x^2 -(7/3)x = 2 does that make sense?

OpenStudy (amistre64):

or do you want to solve this in another way? its up to you

OpenStudy (amistre64):

i gotta know which way you want to take this...they all lead to the same answer :)